# Approximate $2$-hop neighborhoods on incremental graphs: An efficient lazy approach

### v1

@@ 5.2. Results
As we remarked in the introduction, we focused on the case of undirected graphs for ease of exposition and for the sake of space, but our approach extends seamlessly to directed graphs, such as some of the real examples we consider in this section. In such cases, the only caveat to keep in mind is that we define the $h$-ball of a vertex $u$ as the subset of vertices that are reachable from $u$ over a directed path traversing at most $h$ edges. [^9] One could alternatively define the $h$-ball of $u$ as the subset of vertices from which it is possible to reach $u$ traversing at most $h$ directed edges. We stick to the former definition, which is more frequent in social network analysis.
@@
where $i$ refers to the $2$-ball size of the $i$-th sampled vertex. For each dataset, this index is computed for each of $10$ independent runs at each of the $3$ timestamps we consider. Results are summarized for all datasets and combinations $(\varphi,k)$ we consider in Table 2 for the last timestamp ($100\%$ of edge insertions). Results for other timestamps ($50\%$ and $75\%$) are similar and are omitted for the sake of space (see Appendix B for complete results). The main takeaways here are that i) the additional error introduced by the use of sketches (which can be controlled by varying the size of the sketch) is relatively modest; ii) even relatively large values of $\varphi$ and/or small values of $k$ result in performances that are close to those of the baseline that uses sketches to represent balls, but naively performs all light updates. The effect of $k$ is more pronounced when $\varphi$ is large, contributing to a reduction in both error and, to a lesser extent, variance. Finally, we note that although the effect of increasing $k$ can be achieved by decreasing $\varphi$, our analysis in Section 4 suggests that the parameter $k$ provides robustness against worst-case scenarios. Even though the analysis of the impact of sketches on error is out of the scope of this paper, the results in the previous paragraph and the small difference with the baseline (Table 2) when using the sketches further suggest that the analysis of random sequences may also apply to real sequences.
@@
With the same goal as in the previous experiment, we now evaluate the quality of our lazy approach policy combined with the use of sketches in another graph mining task: the Jaccard similarity estimation for $2$-hop neighborhoods. The sketch used to represent the $1$- and $2$- balls is the well-known $h$-minhash signature (broder1997resemblance, ), with $h=100$ hash functions.
@@
Considering all vertex pairs in the entire graph would be computationally prohibitive, as their number is too large. Moreover, many of these pairs would have an extremely low Jaccard similarity, making it difficult to estimate them with a reasonably small error (broder2000identifying, ; broder2001completeness, ). Therefore, to better evaluate our algorithm’s and baseline’s quality, we need vertex pairs with a sufficiently high similarity in their $2$-hop neighborhoods. To address this, we adopted a similar sampling process as before: we selected the $5000$ vertices with the largest $2$-hop neighborhoods at the end of the edge insertion sequence and randomly chose $1000$ pairs whose similarity is at least $0.2$. On that sample, we evaluated the MAPE of the Jaccard similarity estimate computed using the $h$-minhash signatures. Table 3 reports the results at the end of the insertion sequence, while results for the other timestamps ($50\%$ and $75\%$) are similar and are omitted for sake of space (see Appendix B for complete results).
@@
Finally, we measured the running times of our algorithm and the corresponding speed-up with respect to the baseline. In the following, we report and discuss the results for the task of size estimation, using probabilistic counters. The results for the task of Jaccard similarity estimation are analogous, and thus reported in Table 12 in Appendix B.
Table 4 reports the average speed-up of our algorithm with respect to the naive baseline for the same combinations $(\varphi,k)$ considered previously, for all datasets except com-friendster. Speed-ups are computed in terms of total update time, i.e., the total time it takes to process the whole insertion sequence. In Table 5 reports the overall processing times on com-friendster, for a subset of the combinations of $\varphi$ and $k$. It should be noted that the baseline did not complete within a reasonable amount of time in this case. Finally, Figure 6 illustrates the average time cost per operation for the baseline, as well as the slowest and fastest parameter settings of $\varphi$ and $k$ in our algorithm. These experiments clearly highlight that the number of lazy updates is the crucial factor affecting performance and that our approach is very effective at addressing this problem, resulting in considerable speed-ups that grow with the size of the graph. These results regarding processing times, in conjunction with previous findings, underscore the significant advantage in terms of speed at the expense of a modest and acceptable reduction in query quality which, due to the necessary use of sketches, is never totally accurate.
Figure 6. Average time per insertion operation (in seconds). The baseline time for soc-friendster dataset is not reported, since it exceed a time limit of $36$ hrs.
@@
Table 5. Total running time for ball size estimation using probabilistic counters, for com-friendster dataset. The time for the baseline algorithm is not reported since it exceed a time limit of $36$ hrs.
@@ 6. Discussion and outlook
In this work, we showed that relatively simple, lazy update algorithms can play a key role to efficiently support queries over multi-hop vertex neighborhoods on large, incremental graphs. At the same time, this work leaves a number of open questions that might deserve further investigation. A first, obvious direction is extending our approach to handle edge deletions. While our approach is potentially useful in the general case, the main problem here is handling deletions when compact, sketch-based data structures are used to represent $1$- and $2$-balls. Some recent contributions address similar issues in dynamic data streams (BSS20, ; CGPS24, ), but extending our analyses to this general case does not seem straightforward. Another interesting direction is investigating strategies to handle queries over $h$-balls when $h>2$, for example maintaining their sizes under dynamic updates. In this case, each edge addition/deletion potentially has cascading effects over $h$-hops. Optimizing (amortized) update costs in this general setting does not seem trivial and we conjecture that a dependence on $h$ might be necessary. Finally, we remark that our algorithms are inherently local, i.e., whether or not to perform updates involving any vertex $v$ only depends on $v$’s immediate neighborhood. As a consequence, a potentially interesting avenue for further research is to investigate distributed variants of our approach, possibly with massive parallel architectures in mind (lee2012parallel, ).
@@ A.3. Proof of Lemma 4
Appendix B Further experimental results
In this section, we present the comprehensive set of results, providing an exhaustive overview and offering a detailed picture that enriches the information already illustrated in Section 5.
@@ Appendix B Further experimental results
$k$ $\varphi$
$0.1$ $0.25$ $0.5$ $0.75$ $1$
linux 0 6.73x 8.01x 8.67x 8.96x 9.08x
2 6.35x 7.36x 7.85x 8.03x 8.11x
4 6.04x 6.90x 7.26x 7.38x 7.45x
8 5.52x 6.13x 6.35x 6.40x 6.45x
fb-wosn 0 2.55x 3.13x 3.47x 3.63x 3.70x
2 2.38x 2.80x 3.04x 3.13x 3.18x
4 2.27x 2.60x 2.76x 2.81x 2.86x
8 2.07x 2.28x 2.37x 2.38x 2.41x
enron 0 1.85x 1.99x 2.07x 2.10x 2.12x
2 1.77x 1.88x 1.94x 1.96x 1.97x
4 1.70x 1.80x 1.84x 1.86x 1.86x
8 1.59x 1.65x 1.67x 1.68x 1.68x
flickr 0 12.36x 17.15x 20.52x 22.13x 23.10x
2 11.18x 14.69x 16.81x 17.66x 18.19x
4 10.36x 13.11x 14.63x 15.22x 15.59x
8 9.08x 10.90x 11.76x 12.10x 12.32x
youtube 0 41.70x 53.71x 62.06x 66.81x 69.02x
2 38.74x 47.56x 52.71x 55.24x 56.28x
4 36.69x 43.71x 47.37x 49.06x 49.79x
8 33.38x 38.10x 40.17x 40.92x 41.30x
(a) fb-wosn-friends
(b) ia-enron-email-all
(c) comm-linux-kernel-reply

### v2

@@ 5.2. Results
As we remarked in the introduction, we focused on the case of undirected graphs for ease of exposition but our approach extends seamlessly to directed graphs, such as some of the real examples we consider in this section. In such cases, the only caveat to keep in mind is that we define the $h$-ball of a vertex $u$ as the subset of vertices that are reachable from $u$ over a directed path traversing at most $h$ edges. [^14] One could alternatively define the $h$-ball of $u$ as the subset of vertices from which it is possible to reach $u$ traversing at most $h$ directed edges. We stick to the former definition, which is more frequent in social network analysis.
@@
where $i$ refers to the $2$-ball size of the $i$-th sampled vertex. For each dataset, this index is computed for each of $10$ independent runs at each of the $3$ timestamps we consider. Results are summarized for all datasets and combinations $(\varphi,k)$ we consider in Table 2 for the last timestamp ($100\%$ of edge insertions). Results for other timestamps ($50\%$ and $75\%$) are similar and are omitted for the sake of clarity (see Appendix B for complete results). The main takeaways here are that i) the additional error introduced by the use of sketches (which can be controlled by varying the size of the sketch) is relatively modest; ii) even relatively large values of $\varphi$ and/or small values of $k$ result in performances that are close to those of the baseline that uses sketches to represent balls, but naively performs all light updates. The effect of $k$ is more pronounced when $\varphi$ is large, contributing to a reduction in both error and, to a lesser extent, variance. Finally, we note that although the effect of increasing $k$ can be achieved by decreasing $\varphi$, our analysis in Section 4 suggests that the parameter $k$ provides robustness against worst-case scenarios. Even though the analysis of the impact of sketches on error is out of the scope of this paper, the results in the previous paragraph and the small difference with the baseline (Table 2) when using the sketches further suggest that the analysis of random sequences may also apply to real sequences.
@@
With the same goal as in the previous experiment, we now evaluate the quality of our lazy approach policy combined with the use of sketches in another graph mining task: the Jaccard similarity estimation for $2$-hop neighborhoods. The sketch used to represent the $1$- and $2$- balls is the well-known $h$-minhash (broder1997resemblance, ), with $h=100$ hash functions.
@@
Considering all vertex pairs in the entire graph would be computationally prohibitive, as their number is too large. Moreover, many of these pairs would have an extremely low Jaccard similarity, making it difficult to estimate them with a reasonably small error (broder2000identifying, ; broder2001completeness, ). Therefore, to better evaluate our algorithm’s and baseline’s quality, we need vertex pairs with a sufficiently high similarity in their $2$-hop neighborhoods. To address this, we adopted a similar sampling process as before: we selected the $5000$ vertices with the largest $2$-hop neighborhoods at the end of the edge insertion sequence and randomly chose $1000$ pairs whose similarity is at least $0.2$. On that sample, we evaluated the MAPE of the Jaccard similarity estimate computed using the $h$-minhash signatures. Table 3 reports the results at the end of the insertion sequence, while results for the other timestamps ($50\%$ and $75\%$) are similar and are omitted for the sake of clarity (see Appendix B for complete results).
@@
Finally, we measured the running times of our algorithm and the corresponding speed-up with respect to the baseline. In the following, we report and discuss the results for the task of size estimation, using probabilistic counters. The performances for the task of Jaccard similarity estimation are analogous, and thus reported in Table 6 in Appendix B.
Table 4 reports the average speed-up of our algorithm with respect to the naive baseline for the same combinations $(\varphi,k)$ considered previously, for all datasets except com-friendster. Speed-ups are computed in terms of total update time, i.e., the total time it takes to process the whole insertion sequence. Table 5 reports the overall processing times on com-friendster, for a subset of the combinations of $\varphi$ and $k$. It should be noted that the baseline did not complete within a reasonable amount of time in this case. Finally, Figure 6 illustrates the average time cost per operation for the baseline, as well as the slowest and fastest parameter settings of $\varphi$ and $k$ in our algorithm. These experiments clearly highlight that the number of lazy updates is the crucial factor affecting performance and that our approach is very effective at addressing this problem, resulting in considerable speed-ups that grow with the size of the graph. These results regarding processing times, in conjunction with previous findings, underscore the significant advantage in terms of speed at the expense of a modest and acceptable reduction in query quality which, due to the necessary use of sketches, is never totally accurate.
Figure 6. Average time per insertion operation (in seconds) in log-scale. The datasets are arranged on the x-axis in increasing order of number of operations, again in log-scale. The baseline time for soc-friendster dataset is not reported, since it exceeded a time limit of $36$ hrs.
@@
Table 5. Total running time for ball size estimation using probabilistic counters, for com-friendster dataset. The time for the baseline algorithm is not reported since it exceeded a time limit of $36$ hrs.
@@
A quick reference for parameter setting.
The trade-off between accuracy, parameter setting and update costs can be summarized as follows: decreasing $\varphi$ increases the frequency of heavy updates (and thus accuracy), while increasing $k$ increases the number of light updates upon a single edge insertion, again improving accuracy. As our experimental results show, real edge sequences exhibit a behavior that is largely consistent with the random permutation model adopted in the analysis proposed in Section 3 and far from the adversarial sequences analyzed in Section 4. In this perspective, the choice of the parameters can be derived by our theoretical results (see Theorem 2 and Lemma 1). However, in practice, by running our algorithms on a diverse set of real networks for a multiplicity of parameter settings, we identified two robust settings, corresponding to different choices for the trade-off between accuracy and efficiency. The first setting is $\varphi=0.5$ and $k=0$, which yields a very fast solution with a reasonable accuracy. Instead, when the accuracy requirement is more stringent, we suggest setting $\varphi=0.25$ and $k=2$, which usually provides a negligible error with a reasonable speed-up.
Finally, the setting of the parameters for the sketches we used (in particular, their sizes) was simply driven by i) the trade-off established by the theoretical analyses in (trevisan/646978.711822, ) and (broder1997resemblance, ; broder2000identifying, ; CGPS24, ) between accuracy and sketch sizes and ii) the amount of memory we could afford. Ultimately, this resulted in the choices described in Section 5.1, where we used $32$-bit integers.
6. Node Centrality Ranking
In this section, we highlight the effectiveness of our approach in efficiently maintaining a ranking of the vertices of an evolving social network with respect to a popular index of centrality, a key mining task in social network analysis. In this respect, we consider the ranking corresponding to harmonic centrality (rochat2009closeness, ; boldi2014axioms, ), a crucial notion in social network analysis, quantifying the importance or influence of a vertex within a network. It is particularly valuable as a refinement of closeness centrality (freeman2002centrality, ), addressing some of its limitations (rochat2009closeness, ; boldi2014axioms, ).
Harmonic centrality.
Given a graph $G=(V,E)$, the harmonic centrality of a vertex $v\in V$ is defined as
$HC(v)=\sum_{u\in V\setminus\{v\}}\frac{1}{d_{G}(v,u)},$
where $d_{G}(v,u)$ denotes the length (i.e., number of edges) of the shortest path from $v$ to $u$ in $G$, and the value $1/d_{G}(u,v)$ is conventionally taken to be 00 when $u$ is not reachable from $v$. It is easy to see that, equivalently, harmonic centrality can be written as follows:
(7) $HC(v)=\sum_{i=1}^{n}\frac{|L_{i}(v)|}{i},$
where, $L_{i}(v)$ denotes the subset of vertices exactly $i$ hops away from $v$. Computing harmonic centrality for all vertices is expensive already in the static setting, in particular, a cost $\Omega(nm)$, corresponding to the naive approach that performs $n$ BFSs, is unavoidable in certain cases (bergamini2019computing, ). Of course, problems are magnified in the dynamic setting, where the quest for efficient algorithms, even with linear update times, is an active and promising research area (Hanauer22DynamicSurvey, ).
Maintaining harmonic centrality in evolving graphs.
From an application point of view, the primary interest is often in ranking vertices based on their harmonic centrality or maintaining the top-$k$ highest scoring vertices (bisenius2018computing, ; crescenzi2020finding, ; putman2019fast, ), rather than calculating the exact values for all of them. In this perspective our lazy approach affords accurate approximation of the actual ranking of the vertices under incremental updates and with update costs that are typically $O(1)$ on real datasets.
Our approach is straightforward: i) we approximate $HC(v)$ by truncating the summation in (7) to its first two terms; ii) rather than maintaining the sets $L_{1}(v)$ and $L_{2}(v)$, we use probabilistic counters to estimate their cardinalities, a strategy that was first explored by Boldi et al. (boldi2013core, ) in the static setting. More formally, truncated harmonic centrality is defined as
$HC_{2}(v)$ $=\sum_{u\in B_{2}(v)\setminus\{v\}}\frac{1}{d(v,u)}=|L_{1}(v)|+\frac{|L_{2}(v)|}{2}$
$=|B_{1}(v)|-1+\frac{|B_{2}(v)|-|B_{1}(v)|}{2}$
Computation of $HC_{2}(v)$ requires knowledge of the $1$-hop and $2$-hop neighborhoods of $v$ ($B_{1}(v)$ and $B_{2}(v)$ respectively). To this purpose, it is natural to approximate $HC_{2}(v)$ using the approximate sizes of $|B_{1}(v)|$ and $|B_{2}(v)|$ ($|\hat{B}_{1}(v)|$ and $|\hat{B}_{2}(v)|$, respectively), as maintained by our algorithm through probabilistic counters. Therefore, we define the approximate truncated harmonic centrality $\widetilde{HC}_{2}(v)$, an estimator of $HC_{2}(v)$, as follows
(8) $HC_{2}(v)\approx\widetilde{HC}_{2}(v)=|\hat{B}_{1}(v)|-1+\frac{|\hat{B}_{2}(v)|-|\hat{B}_{1}(v)|}{2}.$
In the remainder of this section, we show that vertex ranking obtained using $HC(v)$ and $HC_{2}(v)$ are strongly correlated, justifying our approach to a task that would otherwise be hard to solve in practice for medium or large networks. This shows that our approach enables accurate approximations of important graph properties in dynamic settings, without incurring the overhead of full neighborhood tracking.
Empirical analysis.
We perform two experiments to evaluate the effectiveness of our approximation. In both experiments, we compare three different rankings of the nodes, denoted as $\mathrm{\pi}^{*}$, $\mathrm{\pi}$ and $\widetilde{\mathrm{\pi}}$, each one computed, respectively, according to the following criteria: (i) the exact harmonic centrality, (ii) the 2-truncated harmonic centrality, and (iii) the approximated truncated harmonic centrality that makes use of our lazy approach. We use the same $3$ small-medium datasets (namely comm-linux-kernel-reply, fb-wosn-friends and, ia-enron-email-all) used in the first experiment in Section 5.2. The complexity of computing the exact harmonic centrality makes the computation on larger graphs unfeasible.
In the first experiment (Figure 7), we focus on the top-ranking nodes of $\mathrm{\pi}^{*}$, for different datasets, in different (time) snapshots (w.r.t. real edge insertion sequences). For a given value of $\alpha\in\left[0,100\right]$, we select the set $T_{\alpha}^{*}\subseteq V$ consisting of the top $\alpha\%$ nodes of $\mathrm{\pi}^{*}$. We then compute the rank correlation between $\mathrm{\pi}$ and $\mathrm{\pi}^{*}$, and between $\widetilde{\mathrm{\pi}}$ and $\mathrm{\pi}^{*}$, by restricting the rankings to the nodes in $T_{\alpha}^{*}$. This allows us to assess how well the truncated version and the corresponding approximated one preserve the relative importance of real top central nodes.
The correlations between node rankings is computed using both Spearman’s and Kendall’s $\tau$ rank correlation coefficients, while as discussed in Section 5.2, we set the algorithm parameters to $\varphi=0.25$ and $k=2$, achieving a good trade-off between update time and approximation quality [^15] Concerning the sketch, we use the KMV probabilistic counters (trevisan/646978.711822, ) with size $32$.. All reported results are averaged over $10$ independent runs, and are summarized in Figure 7. Overall, the results show that truncated harmonic centrality ($HC_{2}$) preserves a strong correlation with the exact harmonic centrality ($HC$), confirming its effectiveness as a surrogate metric. More importantly, our approximation $\widetilde{HC}_{2}$ achieves correlation scores that are only marginally lower than those of $HC_{2}$. This indicates that the approximation introduced by our method is sufficiently accurate to preserve the relative ranking, making it a good and lightweight alternative for tracking centrality in incremental networks.
In the second experiment, we evaluate how well our approach is able to recall/retrieve the top-ranked nodes of $\mathrm{\pi}^{*}$. Specifically, for different value of $\alpha\in\{1,5,10\}$, we compute the sets $T_{\alpha}^{*},T_{\alpha},\widetilde{T}_{\alpha}$ of the top $\alpha\%$ nodes of $\mathrm{\pi}^{*}$, $\mathrm{\pi}$, and $\widetilde{\mathrm{\pi}}$, respectively.
We then measure the recall of $T_{\alpha}$ and of $\widetilde{T}_{\alpha}$, defined as the fraction of nodes in $T_{\alpha}^{*}$ that also appear in $T_{\alpha}$ (resp. $\widetilde{T}_{\alpha}$), i.e., $|T_{\alpha}\cap T_{\alpha}^{*}|/|T_{\alpha}^{*}|$ (resp., $|\widetilde{T}_{\alpha}\cap T_{\alpha}^{*}|/|T_{\alpha}^{*}|$). This metric captures how effectively our algorithm identifies the truly central nodes, using only approximate $2$-hop neighborhood information.
The results are summarized in Figure 8, and report the average recall over $10$ independent runs. The reported recall values are consistently high across all datasets and at every point along the edge insertion sequence. This shows that the top nodes identified by $\widetilde{HC}_{2}$ closely match those selected by exact harmonic centrality in evolving networks. Such an accurate recovery of the most central nodes further validates the effectiveness of our algorithm.
Figure 7. Spearman’s and Kendall’s $\tau$ correlation coefficients of the $2$-truncated harmonic centrality $HC_{2}$ and its approximation $\widetilde{HC}_{2}$, with respect to the exact harmonic centrality. Each point represents the correlation of the top $\alpha\%$ vertices with respect to the ordering induced by $\pi^{*}$.
(a) Recall with $HC_{2}$
(b) Recall with $\widetilde{HC}_{2}$
Figure 8. Recall of the top $1\%$, $5\%$, and $10\%$ vertices using the $2$-truncated harmonic centrality $HC_{2}$ (on the left), and the approximate harmonic centrality $\widetilde{HC}_{2}$ (on the right). The x-axis indicate different time snapshot along the edge insertion sequence.
[…]
Final remark.
We finally note that alternative (exact or approximate) approaches not considered in this study, e.g., based on on-demand computing or approximating the quantities of interest at query time, are clearly unfeasible for (continuous) queries of the type we consider, each time requiring $n$ local computations, one for each vertex of the graph.
@@ 7. Discussion and outlook
We showed that relatively simple, lazy update algorithms can play a key role to efficiently support queries over multi-hop vertex neighborhoods on large, incremental graphs. At the same time, this work leaves a number of open questions that might deserve further investigation. A first, obvious direction is extending our approach to handle edge deletions. Actually, a simple strategy to manage a deletion of an edge $(u,v)$ is to recompute the 2-balls of the $O(\deg_{u}+\deg_{v})$ affected vertices via 2-layer BFS visits. The cost of this straightforward approach can be amortized on the entire sequence of edge updates whenever deletions are rare. However, the general case is much more challenging: while our approach is potentially useful, one of the main problems here is handling deletions when compact, sketch-based data structures are used to represent $1$- and $2$-balls. Some recent contributions address similar issues in dynamic data streams (BSS20, ; CGPS24, ), but extending our analyses to this general case does not seem straightforward. Another interesting direction is investigating strategies to handle queries over $h$-balls when $h>2$, for example maintaining their sizes under dynamic updates. In this case, each edge addition/deletion potentially has cascading effects over $h$-hops. Optimizing (amortized) update costs in this general setting does not seem trivial and we conjecture that a dependence on $h$ might be necessary. Finally, we remark that our algorithms are inherently local, i.e., whether or not to perform updates involving any vertex $v$ only depends on $v$’s immediate neighborhood. As a consequence, a potentially interesting avenue for further research is to investigate distributed variants of our approach, possibly with massive parallel architectures in mind (lee2012parallel, ).
@@ A.3. Proof of Lemma 4
Appendix B Additional material for Section 5
In this section, we present the comprehensive set of experimental results, providing an exhaustive overview and offering a detailed picture that completes the material available in Section 5.
(a) fb-wosn-friends
(b) ia-enron-email-all
(c) comm-linux-kernel-reply
$k$ $\varphi$
$0.1$ $0.25$ $0.5$ $0.75$ $1$
linux 0 6.73x 8.01x 8.67x 8.96x 9.08x
2 6.35x 7.36x 7.85x 8.03x 8.11x
4 6.04x 6.90x 7.26x 7.38x 7.45x
8 5.52x 6.13x 6.35x 6.40x 6.45x
fb-wosn 0 2.55x 3.13x 3.47x 3.63x 3.70x
2 2.38x 2.80x 3.04x 3.13x 3.18x
4 2.27x 2.60x 2.76x 2.81x 2.86x
8 2.07x 2.28x 2.37x 2.38x 2.41x
enron 0 1.85x 1.99x 2.07x 2.10x 2.12x
2 1.77x 1.88x 1.94x 1.96x 1.97x
4 1.70x 1.80x 1.84x 1.86x 1.86x
8 1.59x 1.65x 1.67x 1.68x 1.68x
flickr 0 12.36x 17.15x 20.52x 22.13x 23.10x
2 11.18x 14.69x 16.81x 17.66x 18.19x
4 10.36x 13.11x 14.63x 15.22x 15.59x
8 9.08x 10.90x 11.76x 12.10x 12.32x
youtube 0 41.70x 53.71x 62.06x 66.81x 69.02x
2 38.74x 47.56x 52.71x 55.24x 56.28x
4 36.69x 43.71x 47.37x 49.06x 49.79x
8 33.38x 38.10x 40.17x 40.92x 41.30x